OpenStudy (togundeji):

3. Find the geometric mean of the pair of numbers 6 and 10. (1 point) √66 √70 60 2√15

OpenStudy (togundeji):

The answer is 60

zepdrix (zepdrix):

Multiply the `two` numbers together, then take the `second` root. \(\large\rm 6\cdot10=60\) \(\large\rm \sqrt{60}=?\)

OpenStudy (togundeji):

Is it 66

zepdrix (zepdrix):

66 is not one of your options. I'm not sure what you're asking. You're asking if the square root of 60 is equal to 66..?

OpenStudy (togundeji):

3. Find the geometric mean of the pair of numbers 6 and 10. (1 point)

OpenStudy (togundeji):

I don't really understand this either @zepdrix

zepdrix (zepdrix):

To find geometric mean, you multiply the numbers together, and take their square root. \(\large\rm \sqrt{6\cdot10}=\sqrt{60}\) 60 contains a perfect square, which can be removed from the root.

zepdrix (zepdrix):

\[\large\rm \sqrt{60}=\sqrt{4\cdot15}=\sqrt{4}\cdot\sqrt{15}\]

zepdrix (zepdrix):

Square root of 4?

OpenStudy (togundeji):

2 @zepdrix

zepdrix (zepdrix):

\[\large\rm \sqrt{60}=2\cdot\sqrt{15}\]k good.

OpenStudy (togundeji):

So the answer is 60?

rishavraj (rishavraj):

@TOgundeji check out option D :)

OpenStudy (togundeji):

OK @rishavraj Yea I got it, thanks. I was staring at my screen wondering what she ment, but now I got it. The answer is D.